OpenStudy (anonymous):

Evaluate -10 + x ÷ 2(9 – x2 ) for x = -6

OpenStudy (anonymous):

-10+x=-16 (9-x^2)=-27 -16/(2*-27)=8/27

OpenStudy (anonymous):

Is this the problem fraction? \[\frac{x-10}{2 \left(9-x^2\right)} \]

OpenStudy (anonymous):

10 + x ÷ 2(9 – x^2) for x = -6 no I am trying to solve for x